# Carolina Torreblanca

### v1

@@ Introduction
Figure 2 illustrates this sequential classification. It reveals three patterns from our classification. First, quantitative, empirical papers are a minority. They represent 41% of all papers pooled across years. Second, within quantitative work, explanatory papers dominate (81%). Descriptive papers constitute only 18% of empirical quantitative papers. Predictive papers remain rare (240 across the entire period). Third, among explanatory studies, model-based methods remain the majority. Design-based methods account for 29% of explanatory papers, suggesting substantial but still limited methodological transformation.
Averages over two decades may mask important temporal dynamics. Figure 3 shows that the share of empirical quantitative papers has increased steadily, reaching about 48% by 2023. Disaggregation by subfield reveals similar upward trends but persistent differences in baseline levels: In 2023, 27% of International Relations papers use quantitative methods, compared to 57% in Comparative Politics and 84% in American Politics.
@@
For the $30,296$ explanatory empirical quantitative studies, which are the main focus of our analysis in the next section, we classify each paper’s primary research design into more than a dozen categories spanning design-based approaches model-based approaches, and other methods. Design-based approaches include experimental studies—field, survey, and lab experiments—and a range of non-experimental identification strategies. These non-experimental methods include (1) methods based on unconfoundedness and overlap, such as matching and reweighting techniques; (2) natural and quasi-experiments; [^1] These studies use exogenous variation in treatment not controlled by the researcher, such as shocks or sudden institutional rules (e.g., natural disasters, draft lotteries, externally implemented policy changes). They fit within unconfoundedness designs, but researchers often analyze them with regression methods without explicitly stating unconfoundedness or selection-on-observables assumptions. (3) instrumental variables designs; (4) regression discontinuity and regression kink designs; and (5) difference-in-differences designs; [^2] This category includes broadly defined causal panel analysis under parallel trends (Chiu et al. 2023). Its implementations include canonical difference-in-differences, two-way fixed effects models, event studies, longitudinal approaches to staggered treatment adoption, and related variants. All assume some version of the parallel trends assumption. (6) the synthetic control method; and (7) other observational strategies that attempt causal inference by explicitly specifying an assignment mechanism and identifying assumptions such as unconfoundedness, parallel trends, continuity, or exclusion restrictions.
@@
Model-based approaches estimate associations using parametric models without an explicit design component, such as linear regression, generalized linear models, structural equation models, and time-series models including ARMA, GARCH, and vector autoregression. Other methods form a residual category, capturing approaches not covered by the pre-specified design-based or model-based categories. Examples include network analysis, text-as-data methods, agent-based modeling, and purely correlational analyses such as simple bivariate correlations and crosstabs.
Although linear regression can recover causal effects in both experimental and non-experimental settings (Angrist and Pischke 2009; Aronow and Samii 2016; Imbens and Xu 2025), we classify a study as design-based only when it explicitly identifies the source of causal variation, such as randomized treatments or shock-induced variation in natural or quasi-experiments. Many regression-based studies in the corpus do not do this. Instead, they follow what is often called a “kitchen sink regression” approach, either failing to separate the treatment from covariates or adding many covariates in the hope of avoiding omitted-variable bias, without articulating an identification strategy distinct from the model specification. We classify these studies as model-based. In fact, a vast majority of model-based studies in the corpus, 71%, rely on simple regression analysis of this form.
@@
For these explanatory studies, we also code whether the paper states key identification assumptions, defined as explicit statements of the core assumption underlying the design (for example, unconfoundedness, parallel trends, continuity, and exclusion restrictions); merely naming a method is not sufficient. We then record whether the paper makes a causal claim. Explicit claims use terms such as “causes,” “effect,” or “impact.” Implicit claims are coded when the paper frames its contribution in terms of causes and consequences or interprets coefficients as causal, even if it avoids direct causal terminology.
Finally, we code several features of the paper and its transparency practices, including the sample size in the main analysis, whether authors fielded original surveys, whether the paper reports placebo tests, and whether it includes a power analysis. We focus on these practices because they appear across designs and are widely used to assess credibility. These variables allow us to document methodological choices, causal reasoning, and transparency practices across the quantitative political science literature.
An important caveat is that we measure these variable by taking at face value what others have reported in their papers, not by assessing how well they executed a particular research design or whether their reports contain mistakes or misrepresentations of methods. Recent replication studies of political science research have found that, across various designs, there remain many problems with the validity of assumptions, statistical power, and errors in execution (e.g., Stommes, Aronow and Sävje 2023; Lal et al. 2024; Chiu et al. 2025).
@@
Having defined our measures, we next describe how we coded them at scale. We developed an LLM-based coding framework and conducted a structured human validation to generate reliable measures from our corpus of political science articles. We first gathered the raw text of all scraped papers and truncated any article exceeding 50,000 tokens (about 35,000 words) to manage cost and ensure processing efficiency. In practice, this affected only a small number of long review articles, and truncation typically removed appendices rather than main empirical content. We then sent each paper’s text, together with a custom prompt, to OpenAI’s gpt-4o model via the batch API. The System Prompt outlined 19 research dimensions to code, including research design, transparency practices, and causal claims (see Supplementary Materials). Using OpenAI’s structured outputs, we constructed a JSON schema to ensure consistent formatting, and the model coded each article according to this schema. To refine the prompt, we iteratively tested eight versions by qualitatively evaluating outputs for a sample of 116 papers. Each author read a couple dozen papers, noted errors, and proposed revisions. After eight rounds, we finalized the prompt and evaluated it on the five key variables.
We validated five core variables with human coders that anchor our main analysis: (1) whether a paper is an empirical quantitative study, (2) its subfield, (3) its primary research design, (4) whether it states key identification assumptions, and (5) whether it makes a causal claim. Subfield and research design are multi-label variables spanning six and seventeen categories; the remaining variables are binary. Task 1 evaluated whether papers were empirical and quantitative using a random sample of 200 papers. Task 2 assessed the remaining variables using another 200 papers identified by the model as empirical and quantitative, with one paper removed after review. Four research assistants coded each paper in pairs, with disagreements resolved by a third coder in consultation with an author. Although modest relative to the corpus, these samples are sufficient to estimate accuracy with useful precision, and additional spot checks across journals yielded similar patterns. Accuracy was high: Task 1 reached 98% agreement with human coders. In Task 2, accuracy was 83% for subfield, 73% for research design (84% among design-based studies), 78% for identification assumptions (90% among design-based studies), and 82% for causal claims (87% among design-based studies).
@@
To evaluate how far and how deep the credibility revolution has reshaped quantitative empirical research in political science, we present descriptive evidence on the scope and pace of methodological change across the discipline. We examine shifts in research designs, empirical practices, and differences in scholarly influence between studies adopting design-based methods and those using traditional approaches.
@@
How have methodological practices shifted in quantitative explanatory research? If the credibility revolution has taken hold, design-based strategies should increasingly displace model-based approaches. Figure 4(a) traces these aggregate trends from 2003 to 2023. Model-based methods (gray) dominated for most of the period, but design-based approaches (black) rose from 15% to 40%. By 2023, the two approaches accounted for nearly equal shares of explanatory quantitative research, while other methods, a residual category that includes analyses without explicit designs, constitute 18%.
@@
Within this overall shift, subfield differences are relatively modest. As Figure 4(b) shows, American Politics began slightly ahead in 2003 and widened its lead through the mid-2010s, but by 2023 all three major subfields show substantial adoption: American Politics at 52%, Comparative Politics at 38%, and International Relations at 43%.
To understand which specific designs drive this growth, Figure 5 breaks down the composition of design-based work. Survey experiments account for much of the increase and become the single most common design by 2023. Difference-in-differences and matching or reweighting methods based on unconfoundedness also expand, while instrumental variables, field experiments, and regression discontinuity designs maintain smaller but steady shares. Natural/quasi-experiments and synthetic control remain rare.
@@
Next, we examine how each method has evolved over time. Figure 6 traces these trajectories for both experimental and non-experimental approaches. Figure 6(a) reports trends for the three experimental approaches. Survey experiments grow rapidly, rising from about 4% of all explanatory quantitative studies in 2003 to more than 15% by 2023, and constitute roughly 45% of all design-based research by the end of the period. Field experiments remain uncommon but show a modest upward trend. By contrast, lab experiments decline sharply after 2016. This decline likely reflects the increasing availability of online and survey-based experimental platforms, which allow researchers to conduct experiments at lower cost, with larger and more diverse samples, and without the logistical constraints of in-person laboratory settings.
Figure 5: Relative share of design-based studies in all explanatory empirical quantitative studies, by year.
@@
Figure 6(b) reports trends for six non-experimental design-based methods. Difference-in-differences (including its various implementations) grows almost monotonically and reaches more than 6% of all explanatory empirical quantitative papers in 2023. Matching and reweighting methods increase to about 5% around 2012 before plateauing through the end of the period. Instrumental variable designs rise from 3% in 2003 to nearly 5% in 2012, then gradually decline over the past decade. Regression discontinuity designs increase slowly but remain near 2% in 2023, reflecting limited opportunities for such designs. Natural and quasi-experiments remain around 1% of explanatory empirical quantitative studies throughout. Synthetic control appears around 2012 and increases modestly in recent years, but still constitutes less than 1% of explanatory empirical quantitative papers.
@@
Taken together, these patterns show that the credibility revolution has altered methodological practice, but unevenly across designs. Survey experiments account for most of the overall shift, followed by difference-in-differences designs, while other design-based approaches have grown only modestly or remained niche.
@@
We found that design-based methods have risen slowly but steadily over the last two decades, reaching parity with traditional model-based approaches only in 2023. This pattern suggests that credibility-revolution practices have diffused at a modest pace and remain far from universal. One concern, however, is that treating all articles equally may mask differences across publishing venues. More influential journals may adopt emerging methodological standards earlier and thereby shape broader disciplinary expectations. If so, equal weighting could understate both the extent and the speed of change occurring in more widely read higher-impact outlets.
To assess this possibility, Figure 7 plots the share of design-based and model-based studies among explanatory quantitative papers when weighting each article by the SJR score of its publishing journal. The left panel presents trends for the 20 highest impact factor journals; the right panel presents all others. In higher-impact journals, design-based approaches rise more quickly and surpass model-based work several years earlier than in the rest of the field. In contrast, model-based methods remain predominant in lower-impact outlets, although design-based work also increases over time. These patterns indicate that the credibility revolution has diffused unevenly across publishing venues, with faster uptake in journals that play a central role in shaping methodological norms in the discipline.

### v2

@@ Introduction
Figure 2 illustrates this sequential classification. We highlight two important patterns from our classification. First, quantitative, empirical papers are a minority, representing 41% of all papers pooled across years. Second, within quantitative work, explanatory papers dominate (81%). Descriptive papers constitute only 18% of empirical quantitative papers. Predictive papers remain rare (240 across the entire period). Averages over two decades may mask important temporal dynamics. Figure 3 shows that the share of empirical quantitative papers has increased steadily, reaching about 48% by 2023. Disaggregation by subfield reveals similar upward trends but persistent differences in baseline levels: In 2023, 27% of International Relations papers use quantitative methods, compared to 57% in Comparative Politics and 84% in American Politics.
@@
For the $30,296$ explanatory empirical quantitative studies, which are the main focus of our analysis in the next section, we classify each paper’s primary research design into three families: design-based, model-based, and other designs. Design-based approaches tie causal identification to a described source of exogenous or as-if-exogenous variation in treatment assignment (for example, randomized assignment, institutional cutoffs, policy shocks, or plausibly exogenous timing) and articulate the corresponding identification assumptions as claims about the assignment mechanism rather than as implications of the regression specification alone. Specifically, these approaches include experimental studies—field, survey, and lab experiments—and a range of non-experimental identification strategies based on unconfoundedness and overlap, such as matching and reweighting techniques; natural and quasi-experiments; [^3] These studies use exogenous variation in treatment not controlled by the researcher, such as shocks or sudden institutional rules (e.g., natural disasters, draft lotteries, externally implemented policy changes). They fit within unconfoundedness designs, but researchers often analyze them with regression methods without explicitly stating unconfoundedness or selection-on-observables assumptions. instrumental variable designs; regression discontinuity and regression kink designs; difference-in-differences designs; [^4] This category includes broadly defined causal panel analysis under parallel trends (Chiu et al. 2023). Its implementations include canonical difference-in-differences, two-way fixed effects models, event studies, longitudinal approaches to staggered treatment adoption, and related variants. All assume some version of the parallel trends assumption. and the synthetic control method. Design-based methods account for 29% of explanatory papers, suggesting substantial but still limited methodological transformation.
@@
Model-based approaches support causal interpretation primarily through parametric or statistical modeling assumptions, including functional-form restrictions and conditional-independence claims embedded in the specification, and often operationalize causal questions via models such as linear regression, generalized linear models, structural equation models, and time-series models (for example, ARMA, GARCH, and vector autoregression). Other methods capture approaches not covered by these categories, including network analysis, text-as-data methods, agent-based modeling, and purely associational analyses such as bivariate correlations and crosstabs.
Because linear regression can recover causal effects in both experimental and non-experimental settings (Angrist and Pischke 2009; Aronow and Samii 2016; Imbens and Xu 2025), we do not take the use linear regression as diagnostic of model-based research. Only studies that interpret coefficients from covariate-adjusted regressions as causal effects without a stated assignment mechanism or design-based identification argument are coded as model-based. In our corpus, 71%, rely on simple regression analysis of this form.
@@
For explanatory studies, we also code whether the paper states key identification assumptions, defined as explicit statements of the core assumption underlying the design (e.g., unconfoundedness, parallel trends, continuity, or exclusion restrictions). We also code whether quantitative explanatory papers make a causal claim. Explicit claims use terms such as “causes,” “effect,” or “impact.” Implicit claims are coded when the paper frames its contribution in terms of causes and consequences or interprets coefficients as causal, even if it avoids direct causal terminology. Initially, we also aimed to assess whether an explanatory study clearly articulated an interpretable estimand. However, consistent with Lundberg et al. (2021), our pilot study indicated that, aside from papers employing experimental, IV, and RD designs—which typically target average treatment effects (ATE), average treatment effects on the treated (ATT), or local average treatment effects (LATE)—very few studies do so. Consequently, we dropped this measure.
We additionally code several features of the analysis, including sample size and the adoption of research practices intended to enhance credibility. These practices include: (1) whether the paper reports placebo tests, (2) whether it indicates that a power analysis was conducted, and (3) whether the hypotheses or the analysis were pre-registered. We focus on these three because they are commonly used across designs and are intended to reduce different forms of inferential error. Placebo tests and pre-analysis plans help reduce the risk of false positives, respectively by detecting spurious results (Eggers et al. 2024) and by limiting undisclosed analytic flexibility, i.e., p-hacking (Brodeur et al. 2024). Power analyses reduce the likelihood of false negatives by ensuring adequate statistical power to detect meaningful effects, a concern that has grown salient in social science research (Arel-Bundock et al. 2026; Lal et al. 2024). Although increasing sample size is not a research practice per se, we treat it as an indirect indicator of false negative sensitivity: if researchers are more attentive to power, they may design studies with larger samples regardless of whether they conduct formal calculations. Together, these measures allow us to document methodological choices, and causal reasoning across the quantitative political science literature.
Two caveats are in order. First, we measure these variables by taking what is reported in papers at face value. We do not assess how well authors executed a particular research design, whether the identifying assumptions they invoke are warranted in the application or not, or whether the paper contains errors or misrepresentations of methods. Second, our taxonomy is descriptive rather than evaluative or normative: it is not intended to encode any ranking in which design-based studies are treated as inherently more or less credible than model-based studies. For any given paper, the assumptions required for identification in either family can hold or fail in practice. Indeed, recent replication studies of political science research suggest that, across designs, problems with assumption validity, statistical power, and execution errors remain present (e.g., Stommes et al. 2023; Lal et al. 2024; Chiu et al. 2025).
@@
We next describe how we coded them at scale. We developed an LLM-based coding framework and conducted a structured human validation to generate reliable measures from our corpus of political science articles. We first gathered the raw text of all scraped papers and truncated any article exceeding 50,000 tokens (about 35,000 words) to manage cost and ensure processing efficiency. In practice, this affected only a small number of long review articles, and truncation typically removed appendices rather than main empirical content. We then sent each paper’s text, together with a custom prompt, to OpenAI’s gpt-4o model via the batch API. The System Prompt outlined 19 research dimensions to code, including research design, transparency practices, and causal claims (see Supplementary Materials). Using OpenAI’s structured outputs, we constructed a JSON schema to ensure consistent formatting, and the model coded each article according to this schema. To refine the prompt, we iteratively tested eight versions by qualitatively evaluating outputs for a sample of 116 papers. Each author read a couple dozen papers, noted errors, and proposed revisions. After eight rounds, we finalized the prompt and evaluated it on the five key variables.
We validated five core variables with human coders that anchor our main analysis: (1) whether a paper is an empirical quantitative study, (2) its subfield, (3) its primary research design, (4) whether it states key identification assumptions, and (5) whether it makes a causal claim. Subfield and research design are multi-label variables spanning six and seventeen categories; the remaining variables are binary. Task 1 evaluated whether papers were empirical and quantitative using a random sample of 200 papers. Task 2 assessed the remaining variables using another 200 papers identified by the model as empirical and quantitative, with one paper removed after review. Four research assistants coded each paper in pairs, with disagreements resolved by a third coder in consultation with an author. Although modest relative to the corpus, these samples are sufficient to estimate accuracy with useful precision, and additional spot checks across journals yielded similar patterns. Accuracy was high: Task 1 reached 98% agreement with human coders. In Task 2, accuracy was 83% for subfield, 73% for research design (84% among design-based studies), 78% for identification assumptions (90% among design-based studies), and 82% for causal claims (87% among design-based studies). See section A.2 for more details.
@@
To evaluate how quantitative empirical research in political science has changed as a consequence of the credibility movement, we present descriptive evidence on the scope and pace of methodological change across the discipline. We examine shifts in research designs, empirical practices, and differences in scholarly reward between studies adopting design-based methods and those using traditional approaches.
@@
How have methodological practices shifted in quantitative explanatory research? If the credibility revolution has taken hold, design-based strategies should increasingly displace model-based approaches. Figure 4(a) traces these aggregate trends from 2003 to 2023. Model-based methods (gray) dominated for most of the period, but design-based approaches (black) rose from 15% to 40%. By 2023, the two approaches accounted for nearly equal shares of explanatory quantitative research, while other methods, a residual category defined above, constitute 18%.
@@
Within this overall shift, subfield differences are relatively modest. As Figure 4(b) shows, American Politics was an earlier adopter of design-based approaches, but by 2023 all three major subfields show substantial adoption: American Politics at 52%, International Relations at 43%, and Comparative Politics at 38%.
Figure A3(a) in the Supplementary Materials reports overall results excluding survey experiments. Within this subset, the proportion of design-based papers is considerably smaller: 12% of quantitative papers in 2003, rising to 27% in 2023. Conversely, model-based methods remain dominant throughout, accounting for more than half of all quantitative papers in 2023. Figure A3(b) in the Supplementary Materials shows the subfield specific trends when excluding survey experiment. This exclusion reduces the share of design-based papers most starkly for American Politics. Specifically, the difference is of 21 percentage points (pp) in 2023, relative to the proportion when survey experiments are included. Conversely, the difference is slightly smaller for IR in 2023, a 14.5 pp difference that same year and smallest for Comparative Politics with a reduction of only 10.5 pp that year.
@@
Next, we examine which specific designs drive this growth and how each method has evolved over time. Figure 5 traces these trajectories for both experimental and non-experimental approaches. Figure 5(a) reports trends for the three experimental approaches. Survey experiments, which account for much of the increase in design-based methods, grow rapidly, rising from about 4% of all explanatory quantitative studies in 2003 to more than 15% by 2023, and constitute roughly 45% of all design-based research by the end of the period. Field experiments remain uncommon, arguably due to their high cost, but show a modest upward trend. By contrast, lab experiments decline sharply after 2016. This decline likely reflects the increasing availability of online and survey-based experimental platforms, which allow researchers to conduct experiments at lower cost, with larger and more diverse samples, and without the logistical constraints of in-person laboratory settings.
@@
Figure 5(b) reports trends for six non-experimental design-based methods. Difference-in-differences (including its various implementations) grows almost monotonically and reaches more than 6% of all explanatory empirical quantitative papers in 2023. Matching and reweighting methods increased to about 5% around 2012 before plateauing through the end of the period. Instrumental variable designs rose from 3% in 2003 to nearly 5% in 2012, then gradually declined over the past decade. Regression discontinuity designs increase slowly but remain near 2% in 2023, reflecting limited opportunities for such designs. Natural and quasi-experiments remain around 1% of explanatory empirical quantitative studies throughout. Synthetic control methods appeared around 2012 and have increased modestly in recent years, but still constitutes less than 1% of explanatory empirical quantitative papers.
Taken together, these patterns show that the credibility revolution has altered methodological practice, but unevenly across designs. Survey experiments, to the extent that they can be conceptualized as design-based methods, account for most of the overall shift, followed by difference-in-differences designs, while other design-based approaches have grown only modestly or remained niche.
@@
We found that design-based methods have risen slowly but steadily over the last two decades, reaching parity with traditional model-based approaches in 2023. This pattern suggests that credibility-revolution practices have diffused at a modest pace and remain far from universal. One concern, however, is that treating all articles equally may mask differences across publishing venues. More influential journals may adopt emerging methodological standards earlier and thereby shape broader disciplinary expectations. If so, equal weighting could understate both the extent and the speed of change occurring in more widely read higher-impact outlets.